Zeldris:

Jessie is collecting gemstones. At the end of her first week, she has 3 rocks. At the end of her second week, she has 5 rocks. At the end of her third week, she has 7 rocks. Write the steps for calculating the numbers in the sequence that represents the number of gemstones in her collection each week. Then calculate how many rocks she has after 13 weeks.

yodaddy1234:

she will have 27 rocks at the end of 13 weeks i think

Zeldris:

thanks

jhonyy9:

1 week


3 rocks = 1×2 +1 2 w
5 r. = 2×2 +1 3 w
7 r. = 3×2 +1 4 w
= 4×2+1 = 9 r 5 w ................ = 5×2+1 = 11 r 13 w ................ = ? hope so much this above written will help you understanding and calcule it easy !!!